OpenStudy (happykiddo):

The lengths of the sides of a rectangle become 5 times greater than the original measurements. Which statement is true about the rectangle? Its area becomes 5 times greater, and perimeter becomes 15 times greater. Its area becomes 15 times greater, and perimeter becomes 5 times greater. Its area becomes 25 times greater, and perimeter becomes 5 times greater. Its area becomes 5 times greater, and perimeter becomes 25 times greater.

OpenStudy (hoblos):

area of original rectangle =A= l*w perimeter of original rectangle =P = 2l + 2w = 2(l+w) if w'=5w and l'=5l A' = l'*w' = 5l*5w = 25lw = 25A P' = 2l'+ 2w' = 10l + 10w = 10(l+w) = 5P so Its area becomes 25 times greater, and perimeter becomes 5 times greater

OpenStudy (happykiddo):

got it thanks :D

OpenStudy (hoblos):

glad to help XD